Experts claim that 5% of crimes are committed by women. Is there enough evidence to reject the claim if in a sample of 50 crimes, 5 were committed by women? Use 5% significant level.


1
Expert's answer
2021-05-07T09:21:58-0400

"H_0 : p = 0.05 \\

H_1 : p ≠ 0.05"

Critical value at 0.05 level = -1.96, 1.96

Rejection region = reject H0 if z < -1.96 or z >1.96

Sample proportion

"\hat{p} = \frac{5}{50} = 0.1"

Test statistics

"z = \frac{\hat{p}-p}{\sqrt{p(1-p)/n}} \\

= \frac{0.1-0.05}{\sqrt{0.05(1-0.05)/50}} \\

= 1.62"

Since test statistics falls in non-rejection region, we do not have sufficient evidence to reject H0.

We conclude at 0.05 level that we fail to support the claim.